OpenStudy (anonymous):

one winter night the temperature fell 15 degrees between midnight and 5 a.m. By 9 a.m. the temperature had doubled from what it was at 5 a.m. .By noon, it had risen another 10 degrees to 32 degrees. what was the temperature at midnight?

OpenStudy (nathan917):

Let's call the temp at midnight "x," and take this problem one piece of information at a time. "the temperature fell 15 degrees fahrenheit between midnight and 5 am." temp at 5am = (x - 15) "By 9 am the temperature had doubled from what it was at 5am." temp at 9am = 2(x -15) = 2x - 30 "By noon, it had risen another 10 degrees to 32 degress fahrenheit" 32 = 2x - 30 + 10 32 = 2x - 20 52 = 2x 26 = x PROOF: 5am : 26 - 15 = 11 9am : 11 * 2 = 22 noon: 22 + 10 = 32